# CaT-GNN: Enhancing Credit Card Fraud Detection via Causal Temporal Graph Neural Networks

### v1

@@ 4.1 Experimental Setup
we adopt one private $\mathbf{f}$inacial $\mathbf{f}$raud $\mathbf{s}$emi-supervised $\mathbf{d}$ataset Xiang et al. (2023), termed FFSD, with the partially labeled transaction records. Same with the definition in section 2, if a transaction is reported by a cardholder or identified by financial experts as fraudulent, the label $y_{v}$ will be 1; otherwise, $y_{v}$ will be 0. In addition, we also validate on two other public fraud detection datasets $\mathbf{Yelpchi}$ and $\mathbf{Amazon}$. $\mathbf{Yelpchi}$ Rayana and Akoglu (2015) compiles a collection of hotel and restaurant reviews from Yelp, in which nodes represent reviews. And there are three kinds of relationship edges among these reviews. $\mathbf{Amazon}$: The Amazon graph McAuley and Leskovec (2013) comprises reviews of products in the musical instruments category, in which nodes represent users, and the edges are the corresponding three kinds of relationships among reviews. The statistics of the above three datasets are shown in Table 1.
@@
In our experiment, the learning rate $l_{r}$ is set to 0.003, and the batch size batch $N_{batch}$ is established at 256. Moreover, the input dropout ratio $r_{dropout}$ is determined to be 0.2, with the number of attention heads $N_{head}$ set to 4, and the hidden dimension $d$ to 256. We employed the Adam optimizer to train the model over $N_{epoch}=100$ epochs, incorporating an early stopping mechanism to prevent overfitting. In GraphConsis, CARE-GNN, PC-GNN and GTAN, we used the default parameters suggested by the original paper. In Semi-GNN and Player2Vec, We set the learning rate to 0.01. In YelpChi and Amazon datasets, the train, validation, and test ratio are set to be 40%, 20%, and 40% respectively. In the FFSD dataset, we use the first 7 months’ transactions as training data, and the rest as test data. Similar to previous work Liu et al. (2021), we repeat experiments with different random seeds 5 times and we report the average and standard error. Experimental results are statistically significant with $p<0.05$. Cat-GNN and other baselines are all implemented in Pytorch 1.9.0 with Python 3.8. All the experiments are conducted on Ubuntu 18.04.5 LTS server with 1 NVIDIA Tesla V100 GPU, 440 GB RAM.
@@
Table 2: Performance Comparison (in percent ± standard deviation) on YelpChi, Amazon and FFSD datasets across five runs. The best performances are marked with bold font, and the second-to-best are shown underlined.
Dataset YelpChi Amazon FFSD
@@ 4.2 Performance Comparison (RQ1)
In the experiment of credit card fraud detection across three distinct datasets, Cat-GNN showcases superior performance metrics compared to its counterparts. First of all, Cat-GNN achieves the highest AUC in all three datasets, with values of 0.9035, 0.9706, and 0.8281 for YelpChi, Amazon, and FFSD, respectively. This indicates that Cat-GNN consistently outperforms other methods in distinguishing between classes across diverse datasets. Focusing on the F1 Score, which balances the precision $P_{i}$ and recall $R_{i}$, Cat-GNN again tops the charts with scores of 0.7783, 0.9163, and 0.7211 for YelpChi, Amazon, and FFSD. This reflects the model’s robustness in achieving high precision while not compromising on recall, which is essential where both false positives and false negatives carry significant consequences. Finally, Cat-GNN’s superiority extends to the AP metric, with the improvement of at least 6.82%, 2.86%, and 13.10% for YelpChi, Amazon and FFSD respectively.
The comparative performance of Cat-GNN is particularly significant when contrasted with previous methods such as Player2Vec, Semi-GNN, and GraphSAGE. For the Amazon dataset, existing state-of-the-art models, like CARE-GNN, PC-GNN, and GTAN, have already proven effective at capturing the inherent correlations within the data. In this context, the benefits of causal intervention may not be as pronounced, possibly due to the dataset’s simpler local structures and more uniform distribution. However, for the FFSD dataset, our methodology exhibits significant performance improvements. This enhancement is attributed to the complex local structures and the prevalence of unlabeled nodes within the dataset. In such scenarios, causal intervention adeptly learns the inherent attribute connections, thereby boosting the model’s generalization. Additionally, learning mixup weights with a linear layer is more reasonable than weighting with importance scores. Similarly, selecting environment nodes based on proportions is more sensible than choosing a fixed number of environment nodes, and the effect is also slightly better. All in all, This superior performance can be ascribed to the integration of causal theory within the Cat-GNN, enhancing its capacity to comprehend the inherent principles of graph attributes, allowing it to discern complex patterns and interactions that other models are unable to effectively capture.
@@ 4.4 Parameter Sensitivity Analysis (RQ3, RQ4)
In our sensitivity analysis of the environmental ratio as demonstrated in the right of Figure 5, we observed that Cat-GNN’s performance on the Amazon dataset is less affected by variations in the training ratio, with AUC fluctuations not exceeding 2%. Conversely, on the FFSD dataset, as the training ratio increases from 5% to 40%, there is a larger fluctuation in Cat-GNN’s performance. This can be attributed to the characteristics of the dataset or the differences in the distribution of labeled data.
@@ 4.5 Model Efficiency (RQ5)
Dataset YelpChi Amazon FFSD
@@ 6 Conclusion & Future Work
Appendix A Future Work
In this paper, we have pioneered the research into integrating causal reasoning with credit card fraud data, systematically constructing a graph neural network solution that addresses out-of-distribution generalization issues caused by local structure inconsistencies in context. Recognizing the potential of causal theory in the development of AI, our future work will delve into more promising forms of data augmentation. For instance, 1): in the field of spatio-temporal data mining, we plan to explore the application of generative diffusion models Guo et al. [2023] for environmental part enhancement; 2) leverage Language Model Learning Liu et al. [2023]; Fang et al. [2024b] for more sophisticated data descriptions and to guide the augmentation of this textual information; 3) graph pruning: Wang et al. [2023b]; Xia et al. [2023]; Wang et al. [2023c, a]; 4) graph sparification: Wu et al. [2023]; Li et al. [2023]; Wang et al. [2024]; Zhang et al. [2024]; Fang et al. [2024a]; 5): graph explainability Fang et al. [2023a, c, 2022]. Furthermore, the method for identifying causal regions can draw on the OAR metric Fang et al. [2023b].

### v2

@@ 4.1 Experimental Setup
we adopt one open-source $\mathbf{f}$inacial $\mathbf{f}$raud $\mathbf{s}$emi-supervised $\mathbf{d}$ataset Xiang et al. (2023), termed S-FFSD [^2] https://github.com/AI4Risk/antifraud, with the partially labeled transaction records. Same with the definition in section 2, if a transaction is reported by a cardholder or identified by financial experts as fraudulent, the label $y_{v}$ will be 1; otherwise, $y_{v}$ will be 0. In addition, we also validate on two other public fraud detection datasets $\mathbf{Yelpchi}$ and $\mathbf{Amazon}$. $\mathbf{Yelpchi}$ Rayana and Akoglu (2015) compiles a collection of hotel and restaurant reviews from Yelp, in which nodes represent reviews. And there are three kinds of relationship edges among these reviews. $\mathbf{Amazon}$: The Amazon graph McAuley and Leskovec (2013) comprises reviews of products in the musical instruments category, in which nodes represent users, and the edges are the corresponding three kinds of relationships among reviews. The statistics of the above three datasets are shown in Table 1.
@@
In our experiment, the learning rate $l_{r}$ is set to 0.003, and the batch size batch $N_{batch}$ is established at 256. Moreover, the input dropout ratio $r_{dropout}$ is determined to be 0.2, with the number of attention heads $N_{head}$ set to 4, and the hidden dimension $d$ to 256. We employed the Adam optimizer to train the model over $N_{epoch}=100$ epochs, incorporating an early stopping mechanism to prevent overfitting. In GraphConsis, CARE-GNN, PC-GNN and GTAN, we used the default parameters suggested by the original paper. In Semi-GNN and Player2Vec, We set the learning rate to 0.01. In YelpChi and Amazon datasets, the train, validation, and test ratio are set to be 40%, 20%, and 40% respectively. In the S-FFSD dataset, we use the first 7 months’ transactions as training data, and the rest as test data. Similar to previous work Liu et al. (2021), we repeat experiments with different random seeds 5 times and we report the average and standard error. Experimental results are statistically significant with $p<0.05$. Cat-GNN and other baselines are all implemented in Pytorch 1.9.0 with Python 3.8. All the experiments are conducted on Ubuntu 18.04.5 LTS server with 1 NVIDIA Tesla V100 GPU, 440 GB RAM.
@@
Table 2: Performance Comparison (in percent ± standard deviation) on YelpChi, Amazon and S-FFSD datasets across five runs. The best performances are marked with bold font, and the second-to-best are shown underlined.
Dataset YelpChi Amazon S-FFSD
@@ 4.2 Performance Comparison (RQ1)
In the experiment of credit card fraud detection across three distinct datasets, Cat-GNN showcases superior performance metrics compared to its counterparts. First of all, Cat-GNN achieves the highest AUC in all three datasets, with values of 0.9035, 0.9706, and 0.8281 for YelpChi, Amazon, and S-FFSD, respectively. This indicates that Cat-GNN consistently outperforms other methods in distinguishing between classes across diverse datasets. Focusing on the F1 Score, which balances the precision $P_{i}$ and recall $R_{i}$, Cat-GNN again tops the charts with scores of 0.7783, 0.9163, and 0.7211 for YelpChi, Amazon, and S-FFSD. This reflects the model’s robustness in achieving high precision while not compromising on recall, which is essential where both false positives and false negatives carry significant consequences. Finally, Cat-GNN’s superiority extends to the AP metric, with the improvement of at least 6.82%, 2.86%, and 13.10% for YelpChi, Amazon and S-FFSD respectively.
The comparative performance of Cat-GNN is particularly significant when contrasted with previous methods such as Player2Vec, Semi-GNN, and GraphSAGE. For the Amazon dataset, existing state-of-the-art models, like CARE-GNN, PC-GNN, and GTAN, have already proven effective at capturing the inherent correlations within the data. In this context, the benefits of causal intervention may not be as pronounced, possibly due to the dataset’s simpler local structures and more uniform distribution. However, for the S-FFSD dataset, our methodology exhibits significant performance improvements. This enhancement is attributed to the complex local structures and the prevalence of unlabeled nodes within the dataset. In such scenarios, causal intervention adeptly learns the inherent attribute connections, thereby boosting the model’s generalization. Additionally, learning mixup weights with a linear layer is more reasonable than weighting with importance scores. Similarly, selecting environment nodes based on proportions is more sensible than choosing a fixed number of environment nodes, and the effect is also slightly better. All in all, This superior performance can be ascribed to the integration of causal theory within the Cat-GNN, enhancing its capacity to comprehend the inherent principles of graph attributes, allowing it to discern complex patterns and interactions that other models are unable to effectively capture.
@@ 4.4 Parameter Sensitivity Analysis (RQ3, RQ4)
In our sensitivity analysis of the environmental ratio as demonstrated in the right of Figure 5, we observed that Cat-GNN’s performance on the Amazon dataset is less affected by variations in the training ratio, with AUC fluctuations not exceeding 2%. Conversely, on the S-FFSD dataset, as the training ratio increases from 5% to 40%, there is a larger fluctuation in Cat-GNN’s performance. This can be attributed to the characteristics of the dataset or the differences in the distribution of labeled data.
@@ 4.5 Model Efficiency (RQ5)
Dataset YelpChi Amazon S-FFSD